haemoglobin contains 0.25% iron by mass.the molar mass of haemoglobin is 89600.the number of iron atoms present in 1 mole of haemoglobin is?
requires the explaination of the chapter.

Grade:11

2 Answers

Saineel Sanjay Prabhudessai
13 years ago


Hi Sanskriti,

Your solution is as follows:

0.25 % by mass means:

0.25 gms of iron in 100 gms haemoglobin.

therefore in 89600 gms i.e. 1 mole of haemoglobin contains:  0.25 x 89600 = 896 x 0.25 = 224 gms of fe

                                                                                             100

But 1mole of Fe i.e. 56 gms comtains 6.023 x 1023 ( Avogadro number. NA) atoms.

therefore 224 gms of Fe will contain:  224 x 6.023 x 1023  = 2.409 x 1024 atoms.

                                                               56

 

 

With Regards,

Saineel.

Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Percentage of iron=0.25%
100gm of haemoglobin has 0.25gm of iron.
89600 have 89600*.25/100 ​= 224gm
The atomic weight of Fe =56
No. of moles present = 224/56 ​= 4
Therefore 4 moles contains = 4*6.023*10^23
= 2.409*10^24

Thanks and Regards